# Deconstructing What Makes a Good Optimizer for Language Models

### v1

@@ 4.1 Adalayer
To study the behavior of adaptive optimizers like Adam, we begin with describing a layer-wise version of Adam which we refer to as Adalayer. Adam, Adafactor and Adalayer all (approximately) store the diagonal second moment matrix, but with coarser and coarser granularity; for a layer of dimension , Adam explicitly maintains the second moment matrix using parameters in the shape of a matrix. Adafactor stores row and column averages of the second moment matrix which serve as a rank-1 approximation to the second moment matrix. Finally, Adalayer stores a single scalar which is the average of the second moment matrix. We will later consider a generalization of Adalayer where instead of averaging second moment over a layer we will average it over a “block” of parameters which can be a subset of a layer. We note that similar algorithms have been studied before by Ginsburg et al. [2019], Agarwal et al. [2020] but we choose to study this variant since it is a direct analogue of Adam and Adafactor. A simplified version of Adalayer optimizer is given in Algorithm 1, other details such as bias correction are kept same as that for Adam.
In Figure 7 we study the behavior of Adalayer across learning rates. To preserve the correspondence with Adam we fix other hyperparameters to be the same: , and . We find that Adalayer has better performance than SGD333Note that in this section we are setting for all optimizers., but it performs worse than Adam and also lacks Adam’s stability across learning rates. The major difference between Adam and Adalayer is the preconditioning done by Adam within a layer. Intuitively, this preconditioning will have large effects in layers where we expect different weights within a layer to have different gradient scales. The first candidate for such a layer is the last layer, since different tokens have widely different frequencies leading to different gradient scales. To test this hypothesis, we run a corrected version444We note that this reasoning also applies to the first layer, but in our ablations applying this correction the first layer did not make a significant difference. of Adalayer where we treat the set of weights feeding into a logit as a separate block. We henceforth refer to Adalayer with this correction as Adalayer*. This is plotted in Figure 7 and we observe that Adalayer* almost recovers the performance as well as a large fraction of the stability of Adam.
@@
4.2 Most language model parameters can be trained with SGD
@@
To test this, we train the last layer with Adalayer* (fixing a learning rate of ) and the rest of the layers with SGD, both with . In Figure 9 (Left) we show the results while sweeping over SGD learning rates from to . While this improves upon the performance of SGD, we do not recover stability of the Adalayer* and Adam baselines. We trace this instability to LayerNorm blocks: Figure 8 (Left) shows that the effective learning rates for the LayerNorm blocks are much smaller, which suggests that they may destabilize at higher SGD learning rates. To ameliorate this, in Figure 9 (Middle and Right) we add LayerNorm parameters to those being trained with Adalayer* and find that this is sufficient to recover both performance and stability of Adalayer*. For the larger 300m model, we find that this even exceeds the performance of Adalayer*.
@@
In the Appendix, we conduct further experiments investigating these ‘hybrid’ variants of SGD. Another plausible remedy to address the small effective learning rates of the LayerNorm blocks is to simply not train these LayerNorm parameters. In Appendix A, Figure 11, we show that this improves performance and stability relative to SGD but does not recover Adalayer* performance. Furthermore, in Figure 12 we train 150m and 300m models using SGD on the matrix blocks and AdaFactor on the last layer and LayerNorm blocks; we find that performance and stability is comparable, but does not exceed that of Adafactor on the entire network.
@@
Note that a caveat of the above results is that we have introduced an additional hyperparameter— SGD learning rate, which we are sweeping over— while keeping the Adalayer* learning rate in the last layer and layer norm layers fixed. While decoupling the learning rates here is needed (due to SGD’s performant learning rates being orders of magnitude higher than that of Adalayer*), this may be responsible for the observed stability. To address this, we perform the following experiment: we train all the layers with Adalayer* (sweeping over learning rate ) but we stop updating the second moment estimates for all layers except the last layer and LayerNorm blocks after initialization. This implies that these layers are effectively being trained by SGD with a fixed learning rate, though unlike the above results, these learning rates are different for different layers. We implement this by passing 1000 batches to initialized 150m and 300m models to obtain second moment estimates for all layers without letting the model take a gradient step, and then allowing the model to train as normal under the same settings as all of our ablations. As in our previous investigation, we fix other hyperparameters to be the same: , and .
@@
In Figure 10 we show the resulting learning rate sweep for freezing Adalayer* learning rate scales at initialization (with the exception of the last layer and LayerNorm). Surprisingly, we find for the 150m model that we can almost entirely recover the stability and performance of Adalayer*. For the 300m model we also match or exceed the performance of Adalayer*, and even nearly match the peak performance of Adam. Note again that this sweeps learning rate across all network parameters. This provides further evidence for the importance of adaptivity in the last layer and LayerNorm, where in contrast we could used fixed ratios from initialization for all other parameters to recover the performance and stability of Adalayer*. In Appendix B we present additional experiments which include allowing Adalayer* to update for only one of the last layer or LayerNorm, or turning off LayerNorm training. None of these partial modifications fully recovers performance and stability.
@@ 5 Discussion and Limitations
After a comprehensive comparison of a variety of optimizers for language modeling, we have found that many optimizers seem to be roughly equivalent both in terms of optimal performance and hyperparameter stability. Diving deeper, we have shown that most of the benefits of adaptive optimizers arise from their treatment of the last layer and LayerNorm parameters. Of course, there are several limitations to our study including the fact that due to computational constraints we only ablate a few architecture decisions, that we only consider one dimensional hyperparameter sweeps, we fix batch size, and that we limit our study to language modeling with a single dataset. Despite these limitations, we believe that the study sheds new light on the fundamentals of optimization for language modeling and suggests that optimizer choice may not be the optimal point of intervention for increasing efficiency.
@@ Acknowledgments
SK, DM and RZ acknowledges support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. NV, DM and RZ are supported by a Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199.
Appendix A Additional experiments: SGD + adaptive variants (Adalayer*, Adafactor)
@@
In this section, we report additional experiments involving training language models with SGD on a fraction of the models’ parameters and an adaptive optimizer on the remaining parameters. Firstly, given that the effective learning rates of the LayerNorm blocks were observed to be small in Figure 8, it is reasonable to ask whether training the LayerNorm parameters is necessary at all; in Figure 11, we show results for training 150m and 300m models using Adalayer* only on the last layer, using SGD on all other matrix blocks, and turning off training for the LayerNorm parameters. This indeed yields greater stability in comparison to Figure 9 (Left) but does not fully recover the performance of Adalayer*, indicating that training LayerNorm parameters helps with performance, which seems more pronounced in the larger model.
@@
Appendix B Additional experiments: freezing Adalayer learning rate ratios
@@
Appendix C Sophia
In this section, we compare Sophia [Liu et al., 2024] to Signum. Note that Signum is a special case of Sophia, achieved by setting . We find that Sophia does not outperform Signum. No significant change in performance was observed when transferring the hyperparameters suggested by Liu et al. [2024] (eg. , , , weight decay), nor when additionally scaling attention by the inverse of layer index which was used in the original Sophia implementation.

### v2

@@ 3.1 Adalayer
To study the behavior of adaptive optimizers like Adam, we begin with describing a layer-wise version of Adam which we refer to as Adalayer. Adam, Adafactor and Adalayer all (approximately) store the diagonal second moment matrix, but with coarser and coarser granularity; for a layer of dimension , Adam explicitly maintains the second moment matrix using parameters in the shape of a matrix. Adafactor stores row and column averages of the second moment matrix which serve as a rank-1 approximation to the second moment matrix. Finally, Adalayer stores a single scalar which is the average of the second moment matrix. We will later consider a generalization of Adalayer where instead of averaging second moment over a layer we will average it over a “block” of parameters which can be a subset of a layer. We note that similar algorithms have been studied before (Ginsburg et al., 2019; Agarwal et al., 2020) but we choose to study this variant since it is a direct analogue of Adam and Adafactor. A simplified version of Adalayer optimizer is given in Algorithm 1; other details such as bias correction are kept same as that for Adam. In Appendix C, we provide more details about our Adalayer implementation and a correction we make to how Adalayer treats the last layer in order to achieve similar performance and stability to Adam. Specifically, Adalayer when naively applied for each layer is neither performant nor stable to learning rate (Figure 10); however, if we additionally treat the set of weights in the last layer feeding into each logit as its own block, this recovers most of the performance and stability of Adam (see the dotted blue lines in Figure 8). We henceforth refer to Adalayer with this correction as Adalayer*.
@@
3.2 Both the last layer and LayerNorm parameters need adaptivity
@@
To test this, we train the last layer with Adalayer* (fixing a learning rate of ) and the rest of the layers with SGD, both with . In Figure 8 (Left) we show the results while sweeping over SGD learning rates from to . While this improves upon the performance of SGD, we do not recover stability of the Adalayer* and Adam baselines. We trace this instability to LayerNorm blocks: Figure 7 (Left) shows that the effective learning rates for the LayerNorm blocks are much smaller, which suggests that they may destabilize at higher SGD learning rates. To ameliorate this, in Figure 8 (Right) we add LayerNorm parameters to those being trained with Adalayer* and find that this is sufficient to recover both performance and stability of Adalayer*. In Figure 11 in the Appendix, we see this trend continue to hold for 300m and 600m parameter models.
@@
We conduct additional experiments investigating these ‘hybrid’ variants of SGD in the Appendix. Despite most language model parameters being in matrix layers, we show that if we only apply Adalayer* on the matrix layers and train the last layer and LayerNorm layers with SGD, this does not recover performance and stability. Another plausible remedy to address the small effective learning rates of the LayerNorm blocks is to simply not train these LayerNorm parameters. In Appendix 12, Figure 13, we show that this improves performance and stability relative to SGD but does not recover Adalayer* performance. Finally in Appendix E, we replace SGD with AdaSGD (Wang & Wiens, 2020) which still uses a global learning rate but is now adaptive through training; in Figure 16 we show AdaSGD + Adalayer* can recover full Adalayer* performance, but a small gap remains to recover Adam performance.
We acknowledge that while adaptivity for the last layer and LayerNorm parameters seems necessary to retain optimizer performance and stability, there remains a small gap which is fulfilled by adaptivity on the other network parameters; in Figure 14 we also try training 150m and 300m models using SGD on the matrix blocks and Adafactor on the last layer and LayerNorm blocks; we find that performance and stability is comparable, but does not exceed that of Adafactor on the entire network.
@@
Further, note that a caveat of the above results is that we have introduced an additional hyperparameter— SGD learning rate, which we are sweeping over— while keeping the Adalayer* learning rate in the last layer and layer norm layers fixed. While decoupling the learning rates here is needed (due to SGD’s performant learning rates being orders of magnitude higher than that of Adalayer*), this may be responsible for the observed stability. To address this, we perform the following experiment: we train all the layers with Adalayer* (sweeping over learning rate ) but we stop updating the second moment estimates for all layers except the last layer and LayerNorm blocks after initialization. This implies that these layers are effectively being trained by SGD with a fixed learning rate, though unlike the above results, these learning rates are different for different layers. We implement this by passing 1000 batches to initialized 150m and 300m models to obtain second moment estimates for all layers without letting the model take a gradient step, and then allowing the model to train as normal under the same settings as all of our ablations. As in our previous investigation, we fix other hyperparameters to be the same: , and .
@@
In Figure 9 we show the resulting learning rate sweep for freezing Adalayer* learning rate scales at initialization (with the exception of the last layer and LayerNorm). Surprisingly, we find for the 150m model that we can almost entirely recover the stability and performance of Adalayer*. For the 300m model we also match or exceed the performance of Adalayer*, and even nearly match the peak performance of Adam. Note again that this sweeps learning rate across all network parameters. This provides further evidence for the importance of adaptivity in the last layer and LayerNorm, where in contrast we could have used fixed ratios from initialization for all other parameters to recover the performance and stability of Adalayer*. In Appendix F we present additional experiments which include allowing Adalayer* to update for only one of the last layer or LayerNorm, or turning off LayerNorm training. None of these partial modifications fully recover performance and stability.
@@ 4 Discussion and Limitations
After a comprehensive comparison of a variety of optimizers for language modeling, we have found that many optimizers seem to be roughly equivalent both in terms of optimal performance and hyperparameter stability. The only differing axis for practitioners to consider is thus memory constraints, but based on our results, memory considerations for optimizer choice should not affect the performance of their training run, at least across the optimizers tested (i.e. optimizers using diagonal preconditioning). For tuning guidelines, we have reported the optimal learning rates in Appendix B found in our sweeps as a starting point for practitioners in tuning. We also observed that most other parameters are very stable for the optimal learning rate (with exception to momentum) and thus we believe tuning learning rate and momentum should be prioritized if one has a limited compute budget for hyperparameter sweeping.
Diving deeper, we have shown that the treatment of the last layer and LayerNorm parameters is crucial for realizing the benefits of adaptive optimizers. This highlights the following research directions as the most promising for developing optimizers for language model training. Firstly, our results suggest that within the realm of diagonal preconditioning optimizers, there may be room for more memory-efficient alternatives to AdamW to be designed beyond using factorization techniques (eg. Adafactor with momentum and Lion from our initial sweeps). We have provided one direction for cutting down on memory overhead, where we have shown that the need for adaptivity is distributed unevenly across different parameters of the network (with similar ideas already explored in concurrent work (Zhang et al., 2024c)). Secondly, in terms of improving the performance of these models, our work suggests that optimizer choice is unlikely to be a promising approach— if we are limited to diagonal preconditioning optimizers. Thus, it is likely that we need to turn to non-diagonal methods to improve on our current methods; this is indeed where previous work has demonstrated improved performance to Adam (Gupta et al., 2018a).
Of course, there are several limitations to our study including the fact that due to computational constraints we only ablate a few architecture decisions, that we only consider one dimensional hyperparameter sweeps, we fix batch size, and that we limit our study to autoregressive language modeling with a single dataset. Despite these limitations, we believe that the study sheds new light on the fundamentals of optimization for language modeling.
@@ Acknowledgments
SK, DM and RZ acknowledges support from the Office of Naval Research under award N00014-22-1-2377 and the National Science Foundation Grant under award #IIS 2229881. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. NV, DM and RZ are supported by a Simons Investigator Fellowship, NSF grant DMS-2134157, DARPA grant W911NF2010021,and DOE grant DE-SC0022199. RZ and DM are supported by Kempner Institute Graduate Research Fellowships.
Appendix A Related work
One closely related work to ours is Wortsman et al. (2024), which explores the stability of Adam with respect to learning rate. We extend the comparison to other optimizers including SGD, Lion and Adafactor, as well as other hyperparameters including momentum and weight decay.
Optimizers: SGD (Robbins & Monro, 1951) had been the workhorse optimizer for deep learning until 2015, when Adam (Kingma & Ba, 2015) was introduced. Adam is a diagonal preconditioning algorithm that maintains a per-parameter learning rate. Over time, coarser variants of Adam have been proposed, which do not explicitly maintain a learning rate per parameter. Adafactor (Shazeer & Stern, 2018; Zhai et al., 2022) maintains a rank-1 approximation of the preconditioner matrix of Adam. Previous works have also explored Signum (Bernstein et al., 2018; 2019) and have observed its benefits in terms of communication efficiency and fault tolerance. Other works have also explored the similarity of Adam with variants of Signum (Balles & Hennig, 2018), and recently, a close variant of Signum, called Lion (Chen et al., 2023), was discovered using symbolic search over algorithms. Some other optimizers that have recently gained increasing attention from the community include Shampoo (Gupta et al., 2018b) and Sophia (Liu et al., 2024).
Similar to us, prior work has explicitly studied optimizers comparisons across different domains and architectures. For instance, Schmidt et al. (2021) performs a sweep over optimizers, schedulers, and seeds over various vision tasks. As mentioned in the introduction, Kaddour et al. (2024) compares optimizers for masked language modeling and at a fixed model scale. Other works have called to attention better benchmarking practices for optimizers (Schneider et al., 2019; Sivaprasad et al., 2020; Bartz-Beielstein et al., 2020).
Adam and Signum: Many works have explored the relationship between Adam and variants of Signum (Balles & Hennig, 2018; Balles et al., 2020; Kunstner et al., 2023) and empirically demonstrated that Signum (or its close variants) generally performs comparably to Adam. Balles et al. (2020) also argued that signSGD generally performs better when the Hessian is close to diagonal, however, it is unclear if this holds for practical settings. Kunstner et al. (2023) recently demonstrated that Adam and a close variant of Signum exhibit similar performance on a variety of datasets including WikiText-2 (Merity et al., 2017) and SQuAD (Rajpurkar et al., 2016). However, in contrast with our work, all of these are restricted to the setting of vision or masked language modeling, and generally do not sweep over multiple hyperparameters.
Layerwise or blockwise Adam: We study Adalayer, a layerwise version of Adam. This is a special case of the BAGM optimizer (Zheng & Kwok, 2019), specifically BAGM B.1. Similar algorithms have also been studied by previous works (Ginsburg et al., 2019; Agarwal et al., 2020; Liu et al., 2021; Zhang et al., 2024c). In particular, concurrent to our work, Zhang et al. (2024c) propose an algorithm termed Adam-mini, which closely tracks a modified version of Adalayer (called Adalayer*), and demonstrate comparable performance to AdamW. Note that, in our work, Adalayer* is introduced to understand the role played by preconditioning in Adam, and we do not specifically focus on the final performance. Zhang et al. (2024b) empirically study the Hessian spectrum of transformers at initialization and find it to be more heterogeneous across layers as compared to ResNets. They argue that this heterogeneity is evidence towards the importance of Adam in training transformers. In contrast our results (Section 3.2) show that Adam’s preconditioning is particularly important for the last layer and LayerNorm parameters to achieve performance and learning rate stability.
Other related works: For vision transformers, in the fine-tuning phase, Kumar et al. (2024) show that using SGD with frozen embedding parameters leads to competitive performance with Adam. Jelassi et al. (2022) explore the similarity between Adam and normalized gradient descent (Nesterov, 2004) and show that normalized gradient descent on GANs does not suffer from mode collapse, while SGD does. Jiang et al. (2023) empirically demonstrate that Adam steers the parameter trajectory towards better-conditioned regions than SGD. Pan & Li (2022) also show that the parameter trajectory of Adam exhibits much higher directional smoothness than that of SGD. Ahn et al. (2024) show that the performance gap between Adam and SGD exacerbates with depth of the network. In a similar vein to us, Kunstner et al. (2024) show that Adam is less sensitive than gradient descent to class-imbalance present in language tasks; we provide further evidence for the importance of preconditioning the last layer, as well as the LayerNorm parameters.
Appendix B Additional Main Sweep Results
| Optimizer | Optimal Learning Rate |
| Adam | 3.16e-3 (150m), 1e-3 (300m), 1e-3 (600m), 1e-3 (1.2b) |
| Adafactor | 3.16e-3 (150m), 1e-3 (300m), 1e-3 (600m), 1e-3 (1.2b) |
| Lion | 3.16e-4 (150m), 3.16e-4 (300m), 3.16e-4 (600m), 1e-4 (1.2b) |
| Signum | 3.16e-4 (150m), 3.16e-4 (300m), 3.16e-4 (600m), 3.16e-4 (1.2b) |
Appendix C Adalayer
As mentioned in Section 3, to investigate the role of preconditioning on language models for optimizers like Adam, we introduce the Adalayer optimizer for ease of analysis. In this section, we first establish the performance and stability of Adalayer as a reasonable proxy for Adam by making a modification to how Adalayer treats the last layer of the network.
In Figure 10 we study the behavior of Adalayer across learning rates. To preserve the correspondence with Adam we fix other hyperparameters to be the same: , and . We find that Adalayer has better performance than SGD, but it performs worse than Adam and also lacks Adam’s stability across learning rates. The major difference between Adam and Adalayer is the preconditioning done by Adam within a layer. Intuitively, this preconditioning will have large effects in layers where we expect different weights within a layer to have different gradient scales. The first candidate for such a layer is the last layer, since different tokens have widely different frequencies leading to different gradient scales. To test this hypothesis, we run a corrected version444We note that this reasoning also applies to the first layer, but in our ablations applying this correction the first layer did not make a significant difference. of Adalayer where we treat the set of weights feeding into a logit as a separate block. We henceforth refer to Adalayer with this correction as Adalayer*. This is plotted in Figure 10 and we observe that Adalayer* almost recovers the performance as well as a large fraction of the stability of Adam.
Appendix D Additional experiments: SGD + adaptive variants (Adalayer*, Adafactor)
In this section, we report additional experiments involving training language models with SGD on a fraction of the models’ parameters and an adaptive optimizer on the remaining parameters. Firstly, we show that our results from Section 3.2 hold even when training 600m parameter models with Adalayer* applied only on the last layer and LayerNorm parameters in Figure 11.
We also provide further ablations supporting our claim that the largest impact of the adaptivity of Adalayer* is concentrated on the last layer and LayerNorm parameters. Firstly, we train 150m models using Adalayer* on only the matrix parameters, while training the last layer and LayerNorm parameters with SGD. In Figure 12, we see performance improves relative to SGD but we see similar instability at larger learning rates.
@@ Appendix D Additional experiments: SGD + adaptive variants (Adalayer*, Adafactor)
Secondly, given that the effective learning rates of the LayerNorm blocks were observed to be small in Figure 7 (Right), it is reasonable to ask whether training the LayerNorm parameters is necessary at all; in Figure 13, we show results for training 150m and 300m models using Adalayer* only on the last layer, using SGD on all other matrix blocks, and turning off training for the LayerNorm parameters. This indeed yields greater stability in comparison to Figure 8 (Left) but does not fully recover the performance of Adalayer*, indicating that training LayerNorm parameters helps with performance, which seems more pronounced in the larger model.
@@
Appendix E Additional Experiments: AdaSGD + Adalayer*
The previous experiments on SGD + Adalayer* show that although using a fixed learning rate for the matrix parameters largely recovers performance of full Adalayer*, there remains a small gap to full Adalayer* or Adam performance. Can we recover what is remaining? We further investigate the importance of adaptivity at different levels of granularity by training the matrix parameters with AdaSGD (Wang & Wiens, 2020) instead of SGD. AdaSGD still uses a global learning rate but allows this scalar to be adaptive across training— in other words, this is a “global” version of Adalayer* where all matrix parameters are treated as one block.
In Figure 15 we conduct an analogous learning rate sweep for AdaSGD for as we did in Section 2. The performance of AdaSGD is marginally better than SGD, but still suffers from learning rate instability; this is consistent with our previous findings in Figure 7, where we saw that certain parameters had effective learning rates which were multiple orders of magnitude different from other parameters, and thus a single adaptive learning rate across all parameters likely would not suffice for achieving performance and stability.
In Figure 16 we replace SGD with AdaSGD on the matrix parameters and use Adalayer* to train the last layer and LayerNorm parameters. The performance now essentially matches that of full Adalayer* compared to SGD + Adalayer*, showing that the single adaptive learning rate on matrix parameters is largely sufficient to recover Adalayer* performance. However, a gap still remains between Adam performance even with an adaptive learning rate across matrix parameters; it is likely that this small gap is attributed to the need for increased adaptivity granularity for certain matrix parameters.
Appendix F Additional experiments: freezing Adalayer learning rate ratios
@@
Appendix G Sophia
In this section, we compare Sophia (Liu et al., 2024) to Signum. Note that Signum is a special case of Sophia, achieved by setting . We find that Sophia does not outperform Signum. No significant change in performance was observed when transferring the hyperparameters suggested by Liu et al. (2024) (eg. , , , weight decay), nor when additionally scaling attention by the inverse of layer index which was used in the original Sophia implementation.